OpenStudy (anonymous):

5n-2=(n+2) plz solve for n

OpenStudy (campbell_st):

well there is no need to have brackets on the right as you are multiplying by 1 so you have 5n - 2 = n + 2 start by subtracting n from both sides of the equation then add 2 to both sides then solve for n

OpenStudy (anonymous):

n=1

OpenStudy (anonymous):

5n - 2 = n+2 5n=n + 4 4n= 4n n=1